# Tweets to Citations: Unveiling the Impact of Social Media Influencers on AI Research Visibility

### v2

@@ Appendix C Causal Inference
For our causal inference analysis, we draw primarily from the model, techniques, and assumptions presented in Elazar et al. [2023]. In turn, they point interested readers to Imbens and Rubin [2015] and Feder et al. [2021].
@@
In our work, we study the outcome as a binary variable $Y$, indicating if a paper is "highly-cited" ($Y=1$) or "less-cited" ($Y=0$). By introducing our treatment $A$, we can define two cases for any given paper: $Y_{1}$, the outcome if the paper were shared by influencers (when $A=1$), or $Y_{0}$, the outcome if the paper were not shared by influencers (when $A=0$). In actuality, we can only observe exactly one of these for each paper. Rather, the goal of causal inference is to estimate the change in outcome if the treatment were flipped (Figure 12). Specifially, we hope to answer the question, does influencer sharing increase the likelihood of a paper being highly-cited?
@@
To debias the effects of our unobserved confounders, we will utilize the Negative Outcome Control (NOC) framework Card and Krueger [1993], Lipsitch et al. [2010]. Specifically, we will use Difference-in-Difference (DiD) to estimate ATET over two time periods $t\in\{0,1\}$ Sheppard et al. [1999]:
@@
which can be estimated by a fitting a logistic model and bootstrapping for confidence intervals Elazar et al. [2023].
Sofer et al. [2016] establishes that NCO and DiD-based estimators are equivalent, meaning ATET can be rewritten as:
@@
$A$$C$$U$$N$$Y$ptUnobserved Confounders:Creativity, originality, fameptTreatment:InfluencersharingptNegativeControlOutcome:Review scoreptOutcome:Citation countptObserved Confounders:Topic, authors, venue
Figure 12: Causal graph of our setup. $A$ and $Y$ are binary treatment and effect variables, respectively: whether a paper was shared by influencers, and whether the paper is highly cited. As we cannot measure the unobserved confounders (e.g. quality), we estimate the effect of sharing using a negative control outcome variable ($N$). Solid edges represent a directed causal effect, while dashed edges represent an association. Figure and caption modified from Elazar et al. [2023].
@@
The goal of the Negative Outcome Control (NOC) framework Card and Krueger [1993], Lipsitch et al. [2010] is to find a Negative Control Outcome (NCO) variable that is affected by the same confounders $(C,U)$ as the outcome $Y$, but not the treatment. This way, we can de-bias the effects of confounders not captured by the metadata and text embeddings collected from S2. For this, we choose the average conference review score for each paper, collected as described in Section 3. We believe that this sub-sample is representative for the following reasons: (1) influencers often share preprint works before conference acceptance or proceedings are released; (2) our matching accounts for publication venue as a confounder; (3) using bootstrapping on the three tests from Table 5, we find there is not significant evidence to reject that the sub-sample and full dataset citations come from different distributions ($p$-value $>0.05$).
@@
Elazar et al. [2023] use tripartite matching on a number of covariates Zhang et al. [2021], and only sampled from one venue. To control the topic covariate, they use SPECTER embeddings to build 20 topic clusters and minimize distance for numerical variables (author citations, institution rank, etc.)
Instead, we group the treated and control samples based on a mix of exact matching (venue, year, open access) and quantile binning (number of authors, h-index, and author cites). From these groups, we match samples by minimizing the $L_{2}$ difference of text embeddings. This way, we can most accurately capture increased nuance between "viral" topics, techniques, and models.
@@ Appendix D Limitations
Binary Gender Analysis: Our study utilized a gender prediction API that only outputs binary genders (male and female), limiting our ability to capture the full spectrum of gender identities. Consequently, our findings may not accurately represent the diversity of gender identities within the ML community.
Accuracy with Non-Western Names: The gender prediction API may exhibit lower accuracy when analyzing non-Western names. As a result, the study’s conclusions regarding gender disparity might not fully capture the global diversity of our samples.

### v3

@@ Appendix C Causal Inference
For our causal inference analysis, we draw primarily from the model, techniques, and assumptions presented in Elazar et al. (2023). In turn, they point interested readers to Imbens & Rubin (2015) and Feder et al. (2021).
@@
In our work, we study the outcome as a binary variable $Y$, indicating if a paper is ”highly-cited” ($Y=1$) or ”less-cited” ($Y=0$). By introducing our treatment $A$, we can define two cases for any given paper: $Y_{1}$, the outcome if the paper were shared by influencers (when $A=1$), or $Y_{0}$, the outcome if the paper were not shared by influencers (when $A=0$). In actuality, we can only observe exactly one of these for each paper. Rather, the goal of causal inference is to estimate the change in outcome if the treatment were flipped (Figure 12). Specifially, we hope to answer the question, does influencer sharing increase the likelihood of a paper being highly-cited?
@@
To debias the effects of our unobserved confounders, we will utilize the Negative Outcome Control (NOC) framework (Card & Krueger, 1993; Lipsitch et al., 2010). Specifically, we will use Difference-in-Difference (DiD) to estimate ATET over two time periods $t\in\{0,1\}$ (Sheppard et al., 1999):
@@
which can be estimated by a fitting a logistic model and bootstrapping for confidence intervals (Elazar et al., 2023).
Sofer et al. (2016) establishes that NCO and DiD-based estimators are equivalent, meaning ATET can be rewritten as:
@@
$A$$C$$U$$N$$Y$Unobserved Confounders:Creativity, originality, fameTreatment: Influencer sharingNegativeControlOutcome:Review scoreOutcome:Citation countObserved Confounders:Topic, authors, venue
Figure 12: Causal graph of our setup. $A$ and $Y$ are binary treatment and effect variables, respectively: whether a paper was shared by influencers, and whether the paper is highly cited. As we cannot measure the unobserved confounders (e.g. quality), we estimate the effect of sharing using a negative control outcome variable ($N$). Solid edges represent a directed causal effect, while dashed edges represent an association. Figure and caption modified from Elazar et al. (2023).
@@
The goal of the Negative Outcome Control (NOC) framework (Card & Krueger, 1993; Lipsitch et al., 2010) is to find a Negative Control Outcome (NCO) variable that is affected by the same confounders $(C,U)$ as the outcome $Y$, but not the treatment. This way, we can de-bias the effects of confounders not captured by the metadata and text embeddings collected from S2. For this, we choose the average conference review score for each paper, collected as described in Section 3. We believe that this sub-sample is representative for the following reasons: (1) influencers often share preprint works before conference acceptance or proceedings are released; (2) our matching accounts for publication venue as a confounder; (3) using bootstrapping on the three tests from Table 5, we find there is not significant evidence to reject that the sub-sample and full dataset citations come from different distributions ($p$-value $>0.05$).
@@
Elazar et al. (2023) use tripartite matching on a number of covariates (Zhang et al., 2021), and only sampled from one venue. To control the topic covariate, they use SPECTER embeddings to build 20 topic clusters and minimize distance for numerical variables (author citations, institution rank, etc.)
Instead, we group the treated and control samples based on a mix of exact matching (venue, year, open access) and quantile binning (number of authors, h-index, and author cites). From these groups, we match samples by minimizing the $L_{2}$ difference of text embeddings. This way, we can most accurately capture increased nuance between ”viral” topics, techniques, and models.
@@ Appendix D Limitations
Binary Gender Analysis: Our gender prediction utilizes the AMiner Gender Prediction API, which combines the Facebook Generated Names List (Tang et al., 2011) with precision and recall of 95% and 81%, and MagicFG (Gu et al., 2016) with precision and recall of 93% and 94% respectively. Exact numbers for the combined API are not published by the creators but can reasonably be assumed to be higher than each individual method.
Additionally, the gender prediction API only outputs binary genders (male and female), limiting our ability to capture the full spectrum of gender identities. Consequently, our findings may not accurately represent the diversity of gender identities within the ML community.
Accuracy with Non-Western Names: The gender prediction API may exhibit lower accuracy when analyzing non-Western names (Gu et al., 2016). As a result, the study’s conclusions regarding gender disparity might not fully capture the global diversity of our samples.